OpenStudy (anonymous):

log (x+4) + log(x+3)=?

OpenStudy (anonymous):

log (x^2+7x+12)

OpenStudy (ybarrap):

log(x + 4) + log(x + 3) = log(x+4)(x+3) = log(x^2 + 7x + 12) = logx^2 + log7x + log12 = 2*logx + log7 + logx + log12 = (2 + 1)logx + log (7*12) = 3logx + log84 = logx^3 + log84 = log(84*x^3)